At 20.0°C, the vapor pressure of ethanol is 45.0 torr, and the vapor pressure of methanol is 92.0 torr. What is the vapor pressure at 20.0°C of a solution prepared by mixing 24.0  g methanol and 80.0  g ethanol?
Answer #1

According to raoult law

Ptotal= XaPa+XbPb

Mole of methanol= 24 g/32g/mol=0.75

Mole of ethanol= 80/46g/mol=1.74

Total mole = 1.74+0.75= 2.49

Mole fraction of methanol= 0.75/2.49=0.301

Mole fraction of ethanol= 1.74/2.49=0.7

Vapour pressure= (0.301*92)+(0.7*45)= 59.2 torr